OpenStudy (anonymous):

the derivative of 2x^2 I need to know how to solve this using the derivative as a function... (using limits)

OpenStudy (amistre64):

\[\lim_{h->0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

I already know that formula, i need to know how to use it....

OpenStudy (amistre64):

\[\lim\frac{2(x+h)^2-2x^2}{h}\] \[\lim\frac{2(x^2+2xh+h^2)-2x^2}{h}\] \[\lim\frac{2x^2+4xh+2h^2-2x^2}{h}\] \[\lim\frac{\cancel{2x^2}+4xh+2h^2\cancel{-2x^2}}{h}\] \[\lim\frac{4xh}{h}+\frac{2h^2}{h}\] \[\lim\frac{4x}{1}+\frac{2h}{1}\]

OpenStudy (amistre64):

if you know the formula, you know how to use it :/

OpenStudy (amistre64):

now, when h=0; 4x+2h = 4x

OpenStudy (amistre64):

when I first did these the "x" part of the function confused me a bit since they used x as part of the substituions

OpenStudy (amistre64):

f(x+h) should be written better as say: f(a+h) maybe so that you know to replace "x" by whats in the ( )

OpenStudy (anonymous):

ok, now i get it thanks alot

OpenStudy (amistre64):

youre welcome, and good luck :)